OpenStudy (jotopia34):

I can't figure out convergence and divergence tests, I'm going to fail my quiz if I don't figure this out. Please help someone.... Is the series 1/(ln2)^n convergent or divergent and if convergent, to what?

OpenStudy (jotopia34):

\[\sum_{1}^{\infty}\frac{ 1 }{ (\ln2)^n}\]

OpenStudy (phi):

what is ln(2) ?

OpenStudy (jotopia34):

a number

OpenStudy (phi):

roughly 0.69

OpenStudy (jotopia34):

yes, go on

OpenStudy (phi):

what is 0.69 to a big power ?

OpenStudy (jotopia34):

a very large number?

OpenStudy (jotopia34):

so 1/ very large number is very small, to zero?

OpenStudy (phi):

you can do this, if you don't have a calculator 0.7*0.7= 0.49 or about 0.5 0.7*0.7*0.7 is about 0.5*0.7 or 0.35 do you see a pattern ?

OpenStudy (jotopia34):

yes, when you multiply the a1 and a2 you get a3

OpenStudy (phi):

is 0.69^3 bigger or smaller than 0.69^2

OpenStudy (jotopia34):

oops its smaller

OpenStudy (jotopia34):

I am not going to be able to go by that tho cuz we are not allowed calculators

OpenStudy (phi):

and when n gets big you expect 0.69^n to get very small

OpenStudy (jotopia34):

sorry, I got disconnected. I still don't get what you are getting at

OpenStudy (jotopia34):

what type of series is it? Is it geometric? If so, I can't see that in this a_n

OpenStudy (phi):

to do this problem you need to know ln 2 is less than 1 a number between 0 and 1 will get smaller when you raise it to a power

OpenStudy (jotopia34):

does ln2 mean e^2? I don't know what ln really means?

OpenStudy (phi):

ln (short for natural log) means log base e ln 2 is about 0.69 and e^0.69 is 2 but to finish this, 1/tiny number is a big number the terms are getting larger as n gets bigger. converge or diverge ? what do you think ?

OpenStudy (phi):

** I should say e^(ln 2) = 2

OpenStudy (sirm3d):

you can also write \[\frac{1}{(\ln 2)^n}=\left(\frac{1}{\ln 2}\right)^n\]

OpenStudy (phi):

the other way people check for convergence is using tests the ratio test take the ratio of the n+1 term over the nth term if you get a number bigger than 1 you know the sequence is diverging

OpenStudy (jotopia34):

this is insane, It will take me all night to type the equation. What does it converge to?

OpenStudy (phi):

1/0.69 is about 1.44 You could estimate this by saying 1/.7 is 10/7 or 1 3/7 what is 1.44 raised to a big number?

OpenStudy (jotopia34):

But I have to show which test I use and what type of series it is. She never lets us convert ln to numbers.

OpenStudy (phi):

hint: any number bigger than 1 gets bigger when you raise it to a power if n is big enough, you will get huge numbers as n-> infinity, 1.44^n also -> infinity

OpenStudy (jotopia34):

do I just plug in infinity? I thought we weren't supposed to have direct sub. when it makes the exponent infinity,

OpenStudy (phi):

you can use the ratio test however, you will have to know that ln 2 is less than 1 to know that this series diverges.

OpenStudy (jotopia34):

into the limit I meant

OpenStudy (phi):

the -> means approaches as n-> infinity, 1.44^n -> infinity neither actually "get there" but the point is, the terms are getting larger and larger, and the series diverges.

OpenStudy (sirm3d):

try this \[\sum_{n=0}^\infty \frac{1}{(\ln 2)^n}=\sum_{n=0}^\infty \left(\frac{1}{\ln 2}\right)^n\] what kind or series is ^

OpenStudy (jotopia34):

damn thing won't show the equation you typed. I need a minute

OpenStudy (jotopia34):

I thought it weas a geometric series

OpenStudy (sirm3d):

right, it is geometric.

OpenStudy (jotopia34):

but it is not in the geometric form

OpenStudy (phi):

what form do you mean ?

OpenStudy (jotopia34):

ar^n

OpenStudy (phi):

a is 1, r is 1/ln(2) (roughly 1.44)

OpenStudy (jotopia34):

but the n is negative n

OpenStudy (sirm3d):

b^-n is the same as (1/b)^n

OpenStudy (sirm3d):

you are probably looking at (ln 2)^-n you can also write it as (1/ln 2)^n

OpenStudy (jotopia34):

okay, so (1/ln)^-n is same as ln2^n??

OpenStudy (phi):

\[ \frac{1}{(\ln 2)^n} = \left (\frac{1}{ \ln 2} \right)^n\]

OpenStudy (phi):

as sirm already pointed out

OpenStudy (jotopia34):

oh, I see it now...

OpenStudy (sirm3d):

it is geometric, where a = 1 and r = (1/ln 2) as @phi gave

OpenStudy (phi):

you definitely want to put this into a form that gives a positive n

OpenStudy (jotopia34):

ok I see this. Thank you

OpenStudy (jotopia34):

How can I give 2 medals??? grr

OpenStudy (phi):

Can you do the ratio test on this series?

OpenStudy (jotopia34):

if I don't need to, isn't it harder?

OpenStudy (phi):

First, it is good practice Second, you said you have to show which test you used.